# medial quandle and detecting causality

### v3

@@ Abstract
I investigated the capability of medial quandle, quandle whose operation satisfying that $(a_{1}*b_{1})*(a_{2}*b_{2})=(a_{1}*a_{2})*(b_{1}*b_{2})$, to detect causality in (2+1)-dimensional globally hyperbolic spacetime by determining if they can distinguished the connected sum of two Hopf links with an infinite series of relevant three-component links constructed by Allen and Swenberg in 2020, who suggested that any link invariant must be able to distinguish those links for them to detect causality in the given setting. I show that these quandles fail to do so as long as $a\sim b\Leftrightarrow a*b=a$ defines an equivalence relation. The Alexander quandles is an example that this result can apply to.
@@ 1 Introduction
Section 2 of this paper introduced link invariant generated by quandle, an algebraic structure, and specifically for medial quandle. In section 3-4, this paper investigate the capability of medial quandle for distinguishing $L_{2H}$ and Allen-Swenberg links. I show that as long as the relation define on a medial quandle by $a~{}b\Leftrightarrow a*b=a$ is an equivalence relation, then it cannot distinguish $L_{2H}$ and any of Allen-Swenberg link in the series, thus fail to capture causality in the spacetime. Since Alexander quandles satisfy the restriction on qunadle in this result, one application of this result is that all Alexander quandle fail to detect causality in the given spacetime.
@@ 2 Mathematical background
A quandle $Q(*)$ a set $Q$ equipped with a binary operation $*:Q^{2}\to Q~{}|~{}(x,y)\to x*y$ such that the operation satisfy the following requirement: $\forall x,y,z\in Q:$
@@ 2.1 Example: coloring on the trefoil knot
This means any medial quandle with the equivalence relation defined above is unable to distinguish between the any of Allen-Swenberg links and the connected sum of Hopf link by their coloring invariant or the enhanced version, therefore fail to detect causality in the given setting.
@@ 4.2 Lemma 1
Let $Q(*)$ be an arbitrary medial quandle. Let’s first analysis the coloring of $Q(*)$ on the following two structures, called structure ① and ②,respectively. In both structures by labeling 4 of the arcs exposed outside with $x,y,n,m$ representing some element in $Q(*)$, the other 4 out-exposed aces labeled by $x^{\prime},y^{\prime},n^{\prime},m^{\prime}$ can be expressed by a string of quandle operations between $x,y,n,m$. By using the properties of medial quandle given in section 2.6, these expressions can be further simplified to the form of some functions $f_{pq},~{}p,q\in\{x,y,n,m\}$ defined in definition 4.1 acting on one of $x,y,n,m$ The result is shown in figure 4.2 below and the calculation process is shown below.

### v4

# medial quandles’s capability of detecting causality and properties of their coloring on certain links and knots

@@ Abstract
I investigated the capability of medial quandle, quandle whose operation satisfying that $(a_{1}*b_{1})*(a_{2}*b_{2})=(a_{1}*a_{2})*(b_{1}*b_{2})$, to detect causality in (2+1)-dimensional globally hyperbolic spacetime by determining if they can distinguished the connected sum of two Hopf links from an infinite series of relevant three-component links constructed by Allen and Swenberg in 2020, who suggested that any link invariant must be able to distinguish those links for them to detect causality in the given setting. I show that these quandles fail to do so as long as $a\sim b\Leftrightarrow a*b=a$ defines an equivalence relation. The Alexander quandles is an example that this result can apply to. Inspired by this result, I also derived a generalized theorem about the coloring of medial quandles on a specific type of tangles, which help to determine whether this quandles can distinguish between a wider range of knots and links or not.
@@ 1 Introduction
This paper is organized as following: Section 2 introduced link invariant generated by quandle, an algebraic structure, and specifically for medial quandle. In section 3-4, this paper investigate the capability of medial quandle for distinguishing $L_{2H}$ and Allen-Swenberg links. This part eventually show that as long as the relation define on a medial quandle by $a~{}b\Leftrightarrow a*b=a$ is an equivalence relation, then it cannot distinguish $L_{2H}$ and any of Allen-Swenberg link in the series, thus fail to capture causality in the spacetime. Since Alexander quandles satisfy the restriction on qunadle in this result, one application of this result is that all Alexander quandle fail to detect causality in the given spacetime. Inspired by a lemma discovered in section 4, section 5 and 6 define an infinite family of tangles which are then proved to possess an interesting property under the coloring of medial quandle. This property helps to determine medial quandle’s coloring on a wider range of knots and links, and some applications are presented in section 7.
@@ 2 Mathematical background
A quandle $Q(*)$ a set $Q$ equipped with a binary operation $*:(x,y)\to x*y$ such that the operation satisfy the following axioms, $\forall x,y,z\in Q:$
@@ 2.1 Example: coloring on the trefoil knot
This means any coloring invariant or its enhanced polynomial generated by medial quandle with the equivalence relation defined above is unable to distinguish between any of Allen-Swenberg links and the connected sum of Hopf link, therefore this type of medial quandles fail to detect causality in the given setting.
@@ 4.2 Lemma 1
Let $Q(*)$ be an arbitrary medial quandle. Let’s first analysis the coloring of $Q(*)$ on the following two structures, called structure ① and ②,respectively. In both structures by labeling 4 of the arcs exposed outside with $x,y,n,m$ representing some element in $Q(*)$, the other 4 out-exposed aces labeled by $x^{\prime},y^{\prime},n^{\prime},m^{\prime}$ can be expressed by a string of quandle operations between $x,y,n,m$. By using the properties of medial quandle given in section 2.6, these expressions can be further simplified to the form of some functions $f_{pq},~{}p,q\in\{x,y,n,m\}$ defined in definition 4.1 acting on one of $x,y,n,m$ The result is shown in the figure below and the calculation process is shown below.